OpenStudy (anonymous):

Factor the polyomial completely using any method X^6+x^5-X^4-x^3 Medal given for answer

jhonyy9 (jhonyy9):

x5(x+1)-x3(x+1)=(x5-x3)(x+1)

jhonyy9 (jhonyy9):

x5-x3=x3(x2-1)=x3(x+1)(x-1)

OpenStudy (anonymous):

Thanks , medak given

jhonyy9 (jhonyy9):

thank you